OpenStudy (anonymous):

HELP ME FAST WILL GIVE FANS AND MEDALS!!!!!!!!! A cylindrical container holding sugar has a height of 18 inches and a diameter of 4 inches. Which of the following is the closest to the volume of the sugar container? A. 904 cubic inches B. 226 cubic inches C. 113 cubic inches D. 1,017 cubic inches

OpenStudy (anonymous):

FLVS?

OpenStudy (kamibug):

Here's the formula for the volume of a cylinder. :)

OpenStudy (anonymous):

can u help me get it @KamiBug

OpenStudy (kamibug):

Remember that radius is half of the diameter. So the radius is 2. Plug in the height and radius into the formula. V = pi*2^2*18 V = pi*4*18 V = pi*72 V = ? (Remember that pi is 3.14)